lately I have not been to hot on my math tests...One problem i had was..
5(x-1)= 3(x+3)
What I tried: i changed it to 5x-5 = 3x+9

here is where i messed up:i again changed it to 8x-5=9 I thought that I was supposed to combine like terms. Can you please show me the next step in the equation and what the answere should be so I understand?

Answer by Alan3354(69443)

lately I have not been to hot on my math tests...One problem i had was..
5(x-1)= 3(x+3)
What I tried: i changed it to 5x-5 = 3x+9

here is where i messed up:i again changed it to 8x-5=9 I thought that I was supposed to combine like terms. Can you please show me the next step in the equation and what the answere should be so I understand?
=====================
5x-5 = 3x+9
Do the same thing to both sides.
Subtract 3x
2x - 5 = 9
Add 5
2x = 14
Divide by 2
x = 7